5gm of steam at 100c is passed into calorimeter containing liquid.Temperature at liquid rises from 32c to 40c.Then water equivalent of calorimeter and contents is 

Grade:8

1 Answers

Arun
4 years ago

Latent heat of vaporisation of water = 2260kJ/Kg

Specific heat capacity of water = 4185.5 J/Kg

 

Heat lost by steam = heat gained by the water and calorimeter

Formula :

Heat gained by water = mcФ

m = mass of water

c = specific heat capacity of water.

Ф = Change in temperature.  = 40 - 32 = 8

Heat lost by steam = mLv + mcФ

Lv = latent heat of vaporisation

m = mass of steam.

Ф = 100 - 40 = 60

Doing the substitution :

2260000 × 0.005 + 60 × 0.005 × 4185.5 = m × 4185.5 × 8

12555.65 = 33484m

m = 12555.65/33484 = 0.37497 kg

= 0.37497 × 1000 = 374.97 kg

= 374.97kg